what is the fifth root of 3.82?
Found 2 solutions by RAY100, stanbon:
Answer by RAY100(1637)

(3.82)^1/5 = (3.82)^.2 = 1.3074
.
checking,,,1.3074^5 = 3.82,,,,ok

Answer by stanbon(75887)

what is the fifth root of 3.82?
(3.82)^(1/5) = 1.307412620...
Cheers,
Stan H.